OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation y=5x^2+6x-6

OpenStudy (anonymous):

To get x intercept put y=0 in the eq

OpenStudy (anonymous):

y=5x^2+6x-6 0=5x^2+6x-6

hartnn (hartnn):

Compare your quadratic equation with \(y=ax^2+bx+c\) find a,b,c then thex intercepts are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

nw its a quad eq Solve it..)